OpenStudy (anonymous):

the hypotenus if a right triangle is 6 meters long. One leg is one meter longer than the other find the legth of the shorter leg. round to the nearest hundred.

OpenStudy (ash2326):

Let one leg be x , so the other one will be x-1, we have this  trevino do you know Pythagoras theorem??

OpenStudy (anonymous):

yes a^2 + b^2 =c^2

OpenStudy (ash2326):

yeah, now we have \[x^2+(x-1)^2=6^2\] let's expand the brackets \[x^2+x^2-2x+1=36\] we get now \[2x^2-2x-35=0\] Can you solve now??

OpenStudy (anonymous):

so i would now use \[-b \pm \sqrt{b ^{2}-4(a)(c)}\]

OpenStudy (anonymous):

divided by 2a

OpenStudy (ash2326):

yeah use it, you'd get two values of x, find and tell me

OpenStudy (anonymous):

k one sec

OpenStudy (anonymous):

i got 4.713 and -3.713

OpenStudy (ash2326):

you're correct , take the positive value x=4.71 x-1=3.71

OpenStudy (anonymous):

so the answer since we are lookin for the smallest one would be 3.71

OpenStudy (anonymous):

3.713

OpenStudy (ash2326):

no, one leg is 4.713 and the other one is 3.713

OpenStudy (anonymous):

but it wants the legth of the shortest leg so it would be 3.713 right?

OpenStudy (ash2326):

we have to find both the legs, have to solve the triangle

OpenStudy (anonymous):

so c^s - b^2 = a^2

OpenStudy (ash2326):

No, we have c=6 a=x b=x-1 so \[x^2+(x-1)^2=6\] we found x from this quadratic equation. So x and x-1 are the two sides of the right triangle